OpenStudy (abbyrose0105):

Find the solution of four times the square root of the quantity of x plus 2 equals negative 16.

OpenStudy (aaronq):

is it this? \(4\sqrt{x+2}=-16\)

OpenStudy (abbyrose0105):

yes

OpenStudy (abbyrose0105):

I can't type it up like that on my computer haha

OpenStudy (aaronq):

lol you can if you learn \(\LaTeX\) anyway, you wanna do the operations (e.g. multiplication, division) so you isolate the x. \(4\sqrt{x+2}=-16\) divide both sides by 4 \(\sqrt{x+2}=-4\) square both sides \((\sqrt{x+2})^2=(-4)^2\) \(x+2=16\) can you solve it from here?

OpenStudy (abbyrose0105):

Yes thank you so much!

OpenStudy (aaronq):

no problem!

OpenStudy (abbyrose0105):

And this would be an extraneous solution, correct? @aaronq

OpenStudy (aaronq):

sorry i had to look up what that meant. But it is because it doesnt work in the original equation.

OpenStudy (abbyrose0105):

Alright that's what I thought, Thanks again!